OpenStudy (anonymous):

Simplify the sum. State any restrictions on the variables. help with work (x+2)/(x+3) + (10x)/(x^2-9)

OpenStudy (anonymous):

Look to factor out (x^2-9) first because it is a difference of two squares. What does it factor out to?

OpenStudy (anonymous):

You want to form instead of a sum of terms on the left-hand side of the equation, a single ratio (or fraction) of polynomials.

OpenStudy (anonymous):

Therefore looking for a common denominator polynomial is what you want to do.

OpenStudy (yacoub1993):

the problem is that you need both denominators to be the same

OpenStudy (yacoub1993):

Note that you can factor x^2 - 9 into (x + 3)(x - 3).

OpenStudy (yacoub1993):

So we have \[\frac{ x-2 }{ x+3 }+\frac{ 10x }{ \left( x-3 \right) \left( x+3 \right)}\]

OpenStudy (yacoub1993):

Now we are missing a (x -3) on the left fraction so multiply top and bottom by (x -3) and get:

OpenStudy (yacoub1993):

\[\frac{ \left( x-2 \right)\left( x-3 \right) }{ \left( x-3 \right) \left( x+3 \right)}+\frac{ 10x }{\left( x-3 \right) \left( x+3 \right)}\]

OpenStudy (yacoub1993):

So, add the fraction together now that the denominators are the same:

OpenStudy (yacoub1993):

\[\frac{ \left( x-2 \right)\left( x-3 \right)+10x }{ \left( x-3 \right)\left( x+3 \right) }\]

OpenStudy (yacoub1993):

FOIL out the parentheses and simplfy the top and you should be done with that part.

OpenStudy (yacoub1993):

For the restrictions, you need to figure out when the denominator might equal 0, because that is not allowed. You can't divide by 0.

OpenStudy (yacoub1993):

So, (x+3) cannot be zero, which means x cannot be -3. But, x^2 -9 cannot be zero either, which means x^2 cannot be 9 so x cannot be +3 or -3.